A double ionized Lithium atom is hydrogen-like with atomic number 3.

Grade:11

1 Answers

Kevin Nash
9 years ago
(i) En = - 13.6 / n2 Z2 eV / atom
For Li2+ , Z = 3 ∴ En = - 13.6 x 9 / n2 eV / atom
∴ E1 = - 13. 6 x 9 / 1 and E3 = - 13.6 x 9 / 9 = - 13. 6
∆E = E3 – E1 = - 13. 6 – (- 13.6 x 9)
13. 6 x 8 = 108.8 eV /atom
Λ = 12400 / E (in eV) Å = 12400 / 108.8 = 114 Å
(ii) The spectral line observed will be three namely 3 ⟶ 1,
3\rightarrow2, \rightarrow 1.